# Anti de Sitter gauge theory for gravity

### v6

@@ 2.3 Some General Relations
With $\{\theta^{a}\}$ the orthonormal co-frame, the 4-volume form $\epsilon$ is given by the Hodge dual of 111
@@ 3.1 Spinor Representation
[…]
@@
$[\sigma_{ab},\sigma_{cd}]=i[\eta_{ad}\sigma_{bc}+\eta_{bc}\sigma_{ad}-\eta_{ac}\sigma_{bd}-\eta_{bd}\sigma_{ac}]$ (3.7)
@@
$[\sigma_{ab},\gamma_{c}]=i(\gamma_{a}\eta_{bc}-\gamma_{b}\eta_{ac)}$ (3.16)
@@ 3.2 Noether Identities from the Lorentz Symmetry
$i\gamma^{a}(D_{a}-\frac{1}{2}T^{b}\vskip 6.0pt plus 2.0pt minus 2.0pt_{ba})\psi-m\psi=$ 0 0 (3.48)
$i(D_{a}-\frac{1}{2}T^{b}\vskip 6.0pt plus 2.0pt minus 2.0pt_{ba})\bar{\psi}\gamma^{a}+m\bar{\psi}=$ 0 0 (3.49)
@@ 5.1 Gauge Field Lagangian
where $\omega^{ab}=\mu^{ab}$ and $\theta^{c}=l\mu^{4c}$. From (5.4) we have that $[\gamma_{\ast}\mathfrak{h},\mathfrak{t}]\in\mathfrak{t}$, which implies that $\omega$ is the connection one-form of a Lorentz connection on $Q$ and $\theta$ a tensor one-form of type $(AdH,\mathfrak{t})$ on $Q$, where $Ad$ denotes the adjoint representation of the symmetry group. The tensor form $\theta^{a}$ is identified as the canonical form on $O(M)$.
@@
and [,][,][ , ] denotes the exterior product of Lie algebra valued forms. The two-form $\Omega^{ab}$ is the curvature of the Lorentz frame bundle $O(M)$ and $\Theta^{a}$ its torsion two-form. Remark that
@@
$L_{2}(\tilde{\Delta},\ast\tilde{\Delta})$ (5.15)
where $L_{2}$ is an $Ad(G)$-invariant Weil polynomial of degree two on the Lie algebra $\mathfrak{g}$. In the case $\mathfrak{g}=so(2,3)$ here considered, the restriction $\Delta$ of $\tilde{\Delta}$ to $Q$ is a tensor two-form on $Q$. Then the restriction to $Q$ of the Weil form itself, i.e., $L_{2}(\Delta,\ast\Delta)=\gamma^{\ast}L_{2}(\tilde{\Delta},\ast\tilde{\Delta})$, projects to a unique four-form $\mathcal{L}$ on $M$ such that
@@
For matrix groups, an $Ad(G)$-invariant Weil polynomial on of degree two on its Lie-algebra $\mathfrak{g}$ is proportional to
@@
If we put $C=(1/16\pi)(1/g_{l^{2}})$, $g_{l}=l_{p}/l$ where $l_{p}$ is the Planck length, then the Lagrangian defines an Einstein-Cartan theory supplemented with the curvature kinetic energy $-\frac{1}{2}\Omega^{ab}\wedge\ast\Omega_{ab}$ and the torsion kinetic energy $-\frac{1}{2}\Theta^{a}\wedge\ast\Theta_{a}$ One can pull-back the Lagrangian to space-time $M$ by using local sections $\sigma:U\subset M\rightarrow O(M)$. By letting,
@@ 5.3 Conservation Laws
$-\frac{3}{2l^{2}}\epsilon_{abcd}\ast(\Theta^{c}\wedge\theta^{a}\wedge\theta^{d})-\frac{1}{2}\epsilon_{abcd}\ast(\Omega^{ad}\wedge\Theta^{c})-\ast(\Omega^{a}\vskip 6.0pt plus 2.0pt minus 2.0pt_{b}\wedge\ast\Theta_{a})\\ =8\pi l_{p}^{2}\ast DT_{b}^{(M)}-\frac{1}{2}l^{2}\ast DT_{b}^{(\Omega)}+\ast DT_{b}^{(\Theta)}$ (5.41)
@@ 5.4 The Case of Zero Torsion
$D\ast\Omega_{ab}=$ $8\pi g_{l}^{2}S_{ab}$ (5.49)
For obvious reasons, the first equation is called Einstein equation, the second Yang-Mills equation. Equation (5.49) is the field equation for a $SO(1,3)$ gauge theory with $8\pi g_{l}^{2}$ as dimensionless coupling. In component notation, these two equations are,
@@
$D_{d}(R_{ab}\vskip 6.0pt plus 2.0pt minus 2.0pt^{cd})=$ $8\pi g_{l}^{2}S_{ab}\vskip 6.0pt plus 2.0pt minus 2.0pt^{c}$ (5.51)
@@
$T^{a}\vskip 6.0pt plus 2.0pt minus 2.0pt_{b}{}^{(\Omega)}=R^{cdaf}R_{cdbf}-\frac{1}{4}R\delta^{a}\vskip 6.0pt plus 2.0pt minus 2.0pt_{b}R^{cdfg}R_{cdfg}$ (5.52)
@@
Note an important difference wit Einstein-Cartan theory. There, by the second field equation $\varepsilon_{abcd}\Theta^{c}\wedge\theta^{d}=8\pi l_{p}^{2}S_{ab}\;$ the spin tensor must be zero for zero torsion. Here this is no longer the case. The conservation laws (5.45-5.46) reduce to,
$DT_{c}^{(M)}=$ $-\frac{1}{2}\Omega^{ab}(X_{c})\wedge S_{ab}$ (5.53)
$DT_{c}^{(\Omega)}=$ $D\ast\Omega_{ab}\wedge\Omega^{ab}(X_{c})$ (5.54)
So here, only for zero torsion and zero spin current, the matter tensor and curvature kinetic energy are conserved.
From (5.52) we see that $T^{a(\Omega)}_{\;\;b}$ is a symmetric tensor (with zero trace), implying that for zero torsion $T^{a(M)}_{\;\;b}\;$ is also symmetric. This also follows from the Yang-Mills equation (5.49), and the identity $\ast DD\ast\Omega_{ab}=0$ , implying that the left hand side of the Noether identity (5.39) vanishes identically. However, this is also the case for spinless matter, such that
$\theta_{[a}\wedge T_{b]}^{(M)}=0\quad:zero\>torsion\>or\>zero\>spin\>current$ (5.55)
From (5.53-5.54) we we also have the identity,
$DT_{a}^{(M)}=\frac{1}{16\pi g_{l}^{2}}DT_{a}^{(\Omega)}$ (5.56)
@@ 5.5 Solutions
with $\lambda$ arbitrary, is a solution. By (5.57) these are the only solutions. P Baekler and P B Yasskin have proved [23] that the only spherically symmetric solutions of the torsion-free equations (5.57-5.58) are (i) the Schwarzschild-de Sitter metrics with arbitrary cosmological constant $\lambda$; (ii) the Nariai-Bertotti metrics with arbitrary $\lambda$; (iii) a specific family of flat spherical wave metrics. Remark that the cosmological constant $\lambda$ is not a fundamental constant as in Einsteins theory (there is no cosmological term in the Lagrangian in this limit). Rather, $\lambda$ is a property of the solution (an integration constant). Our theory has no matter solutions in the limit here considered. So, $l\rightarrow\infty$ corresponding to an anti-de Sitter $\rightarrow$ Poincaré contraction of the spacetime symmetry group, also corresponds to the ”vacuum limit”. The original Stephenson-Yang theory has serious difficulties with matter solutions and the field equations do not reduce to Poisson’s equation in the Newtonian limit.
@@
For zero torsion the field equations (5.50-5.51) reduce to,
@@
$D_{d}(R_{ab}\vskip 6.0pt plus 2.0pt minus 2.0pt^{cd})=$ $8\pi g_{l}^{2}S_{ab}\vskip 6.0pt plus 2.0pt minus 2.0pt^{c}$ (5.61)
@@
and where $T^{a(M)}_{\;b}$ contains only contributions from ordinary and dark matter. If the curvature $<<1/l_{p}^{-2}$, the Yang-Mills curvature energy-momentum term $T^{a}\vskip 6.0pt plus 2.0pt minus 2.0pt_{b}{}^{(\Omega)}$ can be neglected and the field equation (5.62) reduces to Einstein’s equation.
@@
exists only if also $T^{a(\Omega)}_{\;\;b}=0.$. An explicit calculation of this result is given in appendix A. In this context see also ref.[24].
Inside matter, the Yang-Mills field equation (5.61) imposes restrictions on the solution set of the first field equations. For example for the RW-geometry, this Yang-Mills equation determines the equation of state for the matter present. The only solution found for a RW-geometry with a perfect fluid matter content is the early universe. This is shown in appendix B.
@@ A.1 Geometrical Quantities
$B^{{}^{\prime}}+(B+C)\lambda^{{}^{\prime}}+\dfrac{1}{r}(B-A)=$ 0 0 (A.43)
$D^{{}^{\prime}}+\frac{2}{r}(D-C)=$ 0 0 (A.44)
@@ A.2 Field Equations
$D\ast\Omega_{ab}=$ 0 0 (A.47)
This gives tree Einstein equations: the tt-, rr-, and $\theta\theta=\phi\phi$component,
@@
and two Yang-Mills equations: the tr- and $r\theta=r\phi$component,
$A^{{}^{\prime}}+\frac{2}{r}(A-B)=$ 0 0 (A.51)
$C^{{}^{\prime}}+(B+C)\lambda^{{}^{\prime}}+\frac{1}{r}(C-D))=$ 0 0 (A.52)
@@ A.3.1 Schwarzschild black-hole
We can chose the constant $k$ to be zero by replacing the time coordinate t by another coordinate $te^{-k}$ This is equivalent to replacing $\lambda$ by $\lambda-k$, so that
@@
Finally, from $B-C-D=0\;$ and $\lambda^{{}^{\prime}}=-\mu^{{}^{\prime}}\;$ we have,
@@
Integration with $-2GM\;$ as integration constant gives,
@@ A.3.2 Schwarzschild-de Sitter space
$A+2B+\Lambda=$ 0 0 (A.68)
$A-2C+\Lambda=$ 0 0 (A.69)
$B-C-D+\Lambda=$ 0 0 (A.70)
@@
From $B=-C$ we still have that
@@
Integration with $-2GM\;$ as integration constant gives,
@@ A.3.3 General case
implying $A=-D\;$. So, in this case, the Einstein equations are satisfied for,
@@
Thus $r^{-1}=0\;$ implying $B=C=0\;$ (see definitions (A.12-A.15), in contradiction with (A.84).
@@ B.1 Geometrical Quantities
$D\ast\Omega^{i}\vskip 6.0pt plus 2.0pt minus 2.0pt_{j}=$ 0 0 (B.29)
@@ B.3 Field Equations
In the field equations (5.62) and (5.49), we consider as for the Schwarzschild case, the limit $l\rightarrow l_{p}$ only after solving the field equations and note $\Lambda=\Lambda_{obs}$,
@@
$D\ast\Omega_{ab}=$ 0 0 (B.34)
Using the curvature components and the components of the Yang-Mills energy momentum of the gravitational field, calculated in section B.1, equation (B.33) gives two Einstein equations: the $tt\>$-, and $ii\>$-component,
@@
The $ti\>$-component of the Yang-Mills equation (B.34), gives with (B.28),

### v7

@@ 2.3 Some General Relations
With $\{\theta^{a}\}$ the orthonormal co-frame, the 4-volume form $\epsilon$ is given by the Hodge dual of 11
@@ 3.1 Spinor Representation
Matter fields are represented by tensor $p$-forms $\psi$ of type $(L,E)$ with $L$ denoting the representation of the local Lorentz group in a linear space $E$. Under a local gauge transformation $\Lambda(x)\in SO(1,3)$, a section $\psi(x)$ of $E$ is thus transformed as
@@
$\sigma_{ab},\sigma_{cd}]=i[\eta_{ad}\sigma_{bc}+\eta_{bc}\sigma_{ad}-\eta_{ac}\sigma_{bd}-\eta_{bd}\sigma_{ac}]$ (3.7)
@@
$\sigma_{ab},\gamma_{c}]=i(\gamma_{a}\eta_{bc}-\gamma_{b}\eta_{ac)}$ (3.16)
@@ 3.2 Noether Identities from the Lorentz Symmetry
$i\gamma^{a}(D_{a}-\frac{1}{2}T^{b}\vskip 6.0pt plus 2.0pt minus 2.0pt_{ba})\psi-m\psi=$ $0$ (3.48)
$i(D_{a}-\frac{1}{2}T^{b}\vskip 6.0pt plus 2.0pt minus 2.0pt_{ba})\bar{\psi}\gamma^{a}+m\bar{\psi}=$ $0$ (3.49)
@@ 5.1 Gauge Field Lagangian
where $\omega^{ab}=\mu^{ab}$ and $\theta^{c}=l\mu^{4c}$. From (5.4) we have that $\gamma_{\ast}\mathfrak{h},\mathfrak{t}]\in\mathfrak{t}$, which implies that $\omega$ is the connection one-form of a Lorentz connection on $Q$ and $\theta$ a tensor one-form of type $(AdH,\mathfrak{t})$ on $Q$, where $Ad$ denotes the adjoint representation of the symmetry group. The tensor form $\theta^{a}$ is identified as the canonical form on $O(M)$.
@@
and $[,]$ denotes the exterior product of Lie algebra valued forms. The two-form $\Omega^{ab}$ is the curvature of the Lorentz frame bundle $O(M)$ and $\Theta^{a}$ its torsion two-form. Remark that
@@
[…]
@@
For matrix groups, an $Ad(G)$-invariant Weil polynomial of degree two on its Lie-algebra $\mathfrak{g}$ is proportional to
@@
If we put $C=(1/16\pi)(1/g_{l^{2}})$, $g_{l}=l_{p}/l$ where $l_{p}$ is the Planck length, then the Lagrangian defines an Einstein-Cartan theory supplemented with the curvature kinetic energy $\frac{1}{2}\Omega^{ab}\wedge\ast\Omega_{ab}$ and the torsion kinetic energy $\frac{1}{2}\Theta^{a}\wedge\ast\Theta_{a}$ . One can pull-back the Lagrangian to space-time $M$ by using local sections $\sigma:U\subset M\rightarrow O(M)$. By letting,
@@
Remark that in the Lagrangian (5.23), $\frac{1}{2}\Omega^{ab}\wedge\ast\Omega_{ab}$ is not a topological invariant since we work in a Riemann-Cartan space. It is only after putting torsion to zero that the connection is Levi-Civita, and this term gets proportional to the Euler form.
@@ 5.3 Conservation Laws
$\frac{3}{2l^{2}}\epsilon_{abcd}\ast(\Theta^{c}\wedge\theta^{a}\wedge\theta^{d})-\frac{1}{2}\epsilon_{abcd}\ast(\Omega^{ad}\wedge\Theta^{c})-\ast(\Omega^{a}\vskip 6.0pt plus 2.0pt minus 2.0pt_{b}\wedge\ast\Theta_{a})\\ =8\pi l_{p}^{2}\ast DT_{b}^{(M)}-\frac{1}{2}l^{2}\ast DT_{b}^{(\Omega)}+\ast DT_{b}^{(\Theta)}$ (5.41)
@@ 5.4 The Case of Zero Torsion
$D\ast\Omega_{ab}=$ $0$ (5.49)
For zero torsion one has the Levi-Civita connection, and is $D\ast=\ast D$, such that the Bianchi identity $D\Omega_{ab}=0$ implies $D\ast\Omega_{ab}=0$. From (5.29) we then have that also the spin tensor $S_{ab}$ must be zero. In component notation, these two equations are,
@@
$D_{d}(R_{ab}\vskip 6.0pt plus 2.0pt minus 2.0pt^{cd})=$ $0$ (5.51)
@@
$T^{a}\vskip 6.0pt plus 2.0pt minus 2.0pt_{b}{}^{(\Omega)}=R^{cdaf}R_{cdbf}-\frac{1}{4}\delta^{a}\vskip 6.0pt plus 2.0pt minus 2.0pt_{b}R^{cdfg}R_{cdfg}$ (5.52)
@@
The conservation laws (5.45-5.46) reduce to,
$DT_{a}^{(M)}=$ $0$ (5.53)
$DT_{a}^{(\Omega)}=$ $0$ (5.54)
So, for zero torsion and thus zero spin current, the matter tensor and curvature kinetic energy are conserved.
From (5.52) we see that $T^{a(\Omega)}_{\;\;b}$ is a symmetric tensor (with zero trace), implying that for zero torsion $T^{a(M)}_{\;\;b}\;$ is also symmetric.
@@ 5.5 Solutions
with $\lambda$ arbitrary, is a solution. By (5.55) these are the only solutions. P Baekler and P B Yasskin have proved [23] that the only spherically symmetric solutions of the torsion-free equations (5.55-5.56) are (i) the Schwarzschild-de Sitter metrics with arbitrary cosmological constant $\lambda$; (ii) the Nariai-Bertotti metrics with arbitrary $\lambda$; (iii) a specific family of flat spherical wave metrics. Remark that the cosmological constant $\lambda$ is not a fundamental constant as in Einsteins theory (there is no cosmological term in the Lagrangian in this limit). Rather, $\lambda$ is a property of the solution (an integration constant). Our theory has no matter solutions in the limit here considered. So, $l\rightarrow\infty$ corresponding to an anti-de Sitter $\rightarrow$ Poincaré contraction of the spacetime symmetry group, also corresponds to the ”vacuum limit”. The original Stephenson-Yang theory has serious difficulties with matter solutions and the field equations do not reduce to Poisson’s equation in the Newtonian limit.
@@
The field equations (5.50-5.51) become,
@@
$D_{d}(R_{ab}\vskip 6.0pt plus 2.0pt minus 2.0pt^{cd})=$ $0$ (5.59)
@@
and where $T^{a(M)}_{\;b}$ contains only contributions from ordinary and dark matter. If the curvature $1/l_{p}^{-2}$, the Yang-Mills curvature energy-momentum term $T^{a}\vskip 6.0pt plus 2.0pt minus 2.0pt_{b}{}^{(\Omega)}$ can be neglected and the field equation (5.60) reduces to Einstein’s equation.
@@
exists only if also $T^{a(\Omega)}_{\;\;b}=0$. An explicit calculation of this result is given in App.A. In this context see also ref.[24].
Inside matter, the Yang-Mills field equation (5.61) imposes restrictions on the solution set of the first field equations. For example for the RW-geometry, this Yang-Mills equation determines the equation of state for the matter present. The only solution found for a RW-geometry with a perfect fluid matter content is the early universe. This is shown in App.B.
@@ A.1 Geometrical Quantities
$B^{{}^{\prime}}+(B+C)\lambda^{{}^{\prime}}+\dfrac{1}{r}(B-A)=$ $0$ (A.43)
$D^{{}^{\prime}}+\frac{2}{r}(D-C)=$ $0$ (A.44)
@@ A.2 Field Equations
$D\ast\Omega_{ab}=$ $0$ (A.47)
This gives tree Einstein equations: the tt-, rr-, and $\theta\theta=\phi\phi$ -component,
@@
and two Yang-Mills equations: the tr- and $r\theta=r\phi$ -component,
$A^{{}^{\prime}}+\frac{2}{r}(A-B)=$ $0$ (A.51)
$C^{{}^{\prime}}+(B+C)\lambda^{{}^{\prime}}+\frac{1}{r}(C-D))=$ $0$ (A.52)
@@ A.3.1 Schwarzschild black-hole
We can chose the constant $k$ to be zero by replacing the time coordinate t by another coordinate $te^{-k}$ . This is equivalent to replacing $\lambda$ by $\lambda-k$, so that
@@
Finally, from BCD=0B-C-D=0\; and $\lambda^{{}^{\prime}}=-\mu^{{}^{\prime}}\;$ we have,
@@
Integration with 2GM-2GM\; as integration constant gives,
@@ A.3.2 Schwarzschild-de Sitter space
$A+2B+\Lambda=$ $0$ (A.68)
$A-2C+\Lambda=$ $0$ (A.69)
$B-C-D+\Lambda=$ $0$ (A.70)
@@
From B=CB=-C we still have that
@@
Integration with 2GM-2GM\; as integration constant gives,
@@ A.3.3 General case
implying A=DA=-D\;. So, in this case, the Einstein equations are satisfied for,
@@
Thus $r^{-1}=0\;$ implying B=C=0B=C=0\; (see definitions (A.12-A.15), in contradiction with (A.84).
@@ B.1 Geometrical Quantities
$D\ast\Omega^{i}\vskip 6.0pt plus 2.0pt minus 2.0pt_{j}=$ $0$ (B.29)
@@ B.3 Field Equations
In the field equation (5.60), we cosider as for the Schwarzschild case, the limit $l\rightarrow l_{p}$ only after solving the field equations. We also note $\Lambda=\Lambda_{obs}$,
@@
$D\ast\Omega_{ab}=$ $0$ (B.34)
Using the curvature components and the components of the Yang-Mills energy momentum of the gravitational field, calculated in section B.1, equation (B.33) gives two Einstein equations: the tttt\>-, and iiii\>-component,
@@
The titi\>-component of the Yang-Mills equation (B.34), gives with (B.28),